# Unleashing the Power of Large Language Models for Group POI Recommendations

### v1

@@ Abstract.
Group Point-of-Interest (POI) recommendations aim to predict the next POI that satisfies the diverse preferences of a group of users. This task is more challenging than traditional individual POI recommendations due to complex group decision-making and extremely sparse group-level check-in data. Existing methods for group POI recommendations primarily rely on single ID-based features from check-in data, capturing only statistical correlations and failing to fully utilize the rich semantic information contained in the check-ins, resulting in suboptimal performance. To this end, we propose a framework that unleashes the power of the Large Language Model (LLM) for context-aware group POI recommendations (LLMGPR). Our approach first introduces POI tokens alongside the original word tokens of the LLM, which are initialized by applying the LLM to the rich information of each POI. We then propose a novel sequencing adapter guided by Quantized Low-Rank Adaptation (QLORA) to modify the LLM. The enhanced LLM can learn sequence representations by combining semantic-enhanced POI tokens and rich contextual information including positional encodings and spatio-temporal differences. This approach can be adapted for learning either group or user representations depending on the sequence type. Furthermore, we enhance group representations by aggregating individual member representations with another QLORA-based aggregation adapter and introducing a self-supervised learning task that predicts the purpose of check-in sequences, alleviating the data sparsity issue. Our experimental results demonstrate that LLMGPR outperforms existing methods, effectively addressing group-level data sparsity and providing superior recommendations.
@@
copyright: none
@@ 1. Introduction
With the rapid expansion of Location-Based Social Networks (LBSNs) such as Foursquare and Weeplace, Point-of-Interest (POI) recommendations have become essential for helping individuals discover interesting places. These recommendations are crucial for applications like mobility prediction, route planning, and location-based advertising (Li et al., 2018). While individual POI recommendation systems are effective in addressing personal preferences by leveraging users’ check-in histories, they are limited in scenarios involving group decision-making. As location-based social activities grow, there is a distinct need for recommendations that consider the preferences of multiple users simultaneously, known as Group POI Recommendation. Group-level recommendations are essential in various scenarios, such as friends planning a weekend outing, families on vacation, or colleagues on a business trip, where meeting the collective preferences of all group members is crucial.
@@
Group POI recommendation systems typically involve a two-phase pipeline. In the first phase, the group is treated as a virtual user, allowing traditional individual POI recommendation methods, such as matrix factorization (Lian et al., 2014), transformer-based approaches (Luo et al., 2021), graph-based approaches (Wang et al., 2022), and diffusion-based approaches (Qin et al., 2023), to be directly applied. These methods learn group representations based on group-level check-in data and individual representations based on individual-level check-in data, respectively. In the second phase, these group representations are further enhanced by aggregating the individuals’ representations within the group. Aggregation approaches include the simple average (Vinh Tran et al., 2019), weighted average (Sojahrood and Taleai, 2021), attention mechanism (Sankar et al., 2020), and graph convolution network (Li and Zhu, 2024). Despite their effectiveness, current group POI recommendations typically depend only on the POI-ID sequence data ( including individual and group check-in sequences). This means the models can only capture statistical information from check-ins, while the rich semantic information inherent in the data remains underutilized.
@@
Benefiting from the recent breakthroughs in pre-trained large language models (LLMs), which have significantly advanced the semantic understanding of natural language, some research has started to explore integrating LLMs into POI recommendation systems to better capture the rich semantic information contained in POI data (e.g., POI categories.), thereby alleviating data sparsity and cold-start issues. For example, LLMMove (Feng et al., 2024) and LLM-Mob (Wang et al., 2023) utilize a series of carefully designed prompts to input individual users’ check-in records into pre-trained LLMs, leveraging the commonsense knowledge embedded in the LLMs to better exploit the rich semantic information in check-in data. Additionally, LLM4POI (Li et al., 2024) further fine-tunes the pre-trained LLM using individual check-in data to enhance its performance.
However, current LLM-based POI recommendation methods are primarily designed for individual POI recommendations, making them non-trivial to work for the more challenging group POI recommendations due to the following challenges: (1) Diverse Preferences and Complex Group Decision-making: Unlike individual POI recommendations that only need to satisfy the preferences of a single user, as shown in Figure 1, group POI recommendations must find a consensus among group members on a POI to visit. However, the interests and preferences of individuals within a group are often diverse, and sometimes a group’s preferences are even different from the preferences of each group member (Guo et al., 2021), making it challenging to model the complex process of group decision-making. Existing group POI recommendation works (Liu et al., 2024; Li et al., 2020) are mainly based on the POI-ID sequence data, failing to capture the complex factors influencing group decision-making. On the other hand, existing LLM-based individual POI recommendation methods are typically designed to handle the preferences of a single user, leveraging personal check-in records and profile data. These methods do not account for the diverse individual preferences in a group and the complex group decision-making, which is essential for effective group recommendations. (2) Extreme Data Sparsity: Unlike individual check-ins, which are recorded by single users visiting a POI, group check-in data involves multiple users checking in to the same location together, such as a group of friends checking in together at a restaurant. Group check-ins happen less frequently because they require coordination among several users, making group check-in data significantly sparser than individual check-in data. For example, as discussed in Section 4.1, the group check-in data in two public datasets is approximately 50-100 times sparser than individual check-in data. Therefore, despite the fact that existing LLM-based POI recommendation methods have shown potential in mitigating individual data sparsity by utilizing the semantic information in check-in data, they face greater challenges when dealing with group check-in data. The extremely limited amount of group check-ins makes it difficult to fine-tune models using the sparse group check-in records.
@@
In this work, we introduce a pre-trained large language model-based framework for group POI recommendations (LLMGPR). Specifically, to address the challenge of Diverse Preferences and Complex Group Decision-making, our framework first introduces POI tokens alongside the original word tokens of the LLM. POI embeddings are initialized by applying the LLM to the rich information of each POI (e.g., category and address). By combining a semantic-enhanced POI token and rich contextual information (i.e., positional encodings and spatio-temporal differences) to model each check-in activity, the LLM can capture the complex and diverse factors for the group decision-making beyond the aggregation of individual preferences, such as sequential effect, geographical influence, temporal effect, and semantic effect (Xie et al., 2016). Then, based on the representation of each check-in activity, we propose a novel sequencing adapter under the guidance of Quantized Low-Rank Adaptation (QLORA) to learn the representation of a check-in sequence, which can be used as either group representation or user representation depending on the check-in sequence types.
@@
To address the Extreme Data Sparsity challenge, we first propose a novel aggregation adapter to aggregate individual member representations obtained from their check-in sequences to enhance our group representations. Then, we introduce a novel self-supervised learning (SSL) task to further enhance group representations by learning and predicting sequence purposes (e.g., business trips and family vacations). Specifically, we first designed a prompt that leverages the spatio-temporal and contextual information within each sequence, enabling a pre-trained large language model to assign a category reflecting its underlying purpose, which then serves as the ground truth for self-supervised learning. In this way, the SSL signal helps the LLM capture and utilize additional semantic information of sparse group-level check-in data to refine the group representations. By jointly optimizing the self-supervised learning task (i.e., pretext task) and the adapters for the recommendation task, our framework effectively leverages the rich semantic information in check-in data to provide accurate and robust group-level POI recommendations. In conclusion, the contributions of this work are summarized as follows:
@@
  • We propose LLMGPR, a pre-trained LLM-based framework for the next group POI recommendation. To the best of our knowledge, this is the first work to integrate pre-trained LLMs in the context of group POI recommendations.
@@
  • We introduce semantic-enhanced POI tokens and integrate rich contextual information to represent each group check-in activity, accounting for the complex and diverse factors influencing group decisions. Based on the representation of each check-in activity, we propose a novel sequencing adapter based on QLORA to learn the group representation from the group’s check-in sequence.
  • To address data sparsity of group-level check-in data, we propose two methods to enhance the group representation: we first propose a novel aggregation adapter to combine individual member representations and then introduce a novel self-supervised learning (SSL) task that predicts sequence purposes, empowering the model to obtain more relevant semantic information.
@@ 3. Methodology
3.1. Learning Group Representation
@@
Before establishing the sequence representation semantically, we first introduce POI embeddings. While semantic representations capture the nuanced meanings of POI names and categories, relying solely on semantics may not fully leverage the unique identifiers of POIs. By incorporating POI embeddings through their IDs, we can ensure that each location is distinctly recognized and appropriately contextualized. This method allows us to directly reference specific POIs, enhancing the precision of recommendations and maintaining the integrity of location-specific data. Combining semantic information with POI embeddings provides a comprehensive approach that balances the depth of semantic understanding with the specificity of location identifiers. Formally, we use $\textbf{E}_{poi}\in\mathbb{R}^{P\times d}$ to denote the embedded representations of all POIs, where $P$ is the number of all POIs and $d$ is the embedding size which is the same as the size of the token embedding. Thus, all POIs can be regarded as additional tokens to the large language model. To maintain semantic information, instead of random initialization, the embedding of each POI is initialized by the large language model with a carefully designed prompt which is shown in Figure 3. Given the prompt, we directly use the output of the final transformer block as the initialized POI embedding without going through the linear and softmax layers. In addition, these embeddings will be further refined, making them adaptable for the task of group POI recommendations.
Figure 3. The prompt for initializing POI embeddings.
Given all POI embeddings, we design another prompt shown in Figure 4 to generate the representation of the check-in sequence. When this prompt is processed by the word embedding layer, the respective embedding in $\textbf{E}_{poi}$ will be utilized as the representation of the “POIID”, instead of treating it as a standard word. Similar to the initialization of POI embedding, the output of the final transformer block will be regarded as the embedded representation of the check-in sequence. This process is denoted as:
(4) $\textbf{e}_{g_{n}}=\Theta(\mathcal{X}_{g_{n}},\textbf{E}_{poi}),$
@@
Figure 4. The prompt for generating sequence embeddings.
@@
3.2. Enhancing Group Representations
@@
Figure 5. The prompt for classifying the check-in sequence.
3.3. Self-supervised Leaning Task
As mentioned above, the LLM will be fine-tuned for check-in sequence representation generation and individual preference aggregation. However, the sparsity of user-POI interactions has been a long-standing issue in POI recommendation, which deteriorates further in group settings. To this end, we introduce a self-supervised signal to enhance representations for check-in sequences with the corresponding purposes, such as business trips and tourism, as the travel purposes provide a strong signal to predict the next POI a group/user will visit. Unfortunately, check-in data typically lacks explicit labels indicating the purpose of each check-in sequence. To overcome this limitation, we leverage the powerful capabilities of the LLM to classify each check-in sequence. The LLM analyzes the spatio-temporal and contextual information within the sequence and assigns a category that reflects its underlying purpose. To achieve this, we design the corresponding prompt which is shown in Figure 5, and we have $11$ purpose labels: {Business Trip, Work Commute, Tourism, Short Vacation, Cultural & Entertainment, Shopping & Dining, Social Events, Fitness & Wellness, Daily Commute, Medical & Health Visits, Education & Training}. The details of these labels are provided in Appendix A.1.
Beyond that, those purpose labels are only valid for short sequences, since a purposeful trip won’t last long. Therefore, we divide the whole sequence into subsequences if the time difference between any two consecutive check-in activities is larger than five days. Here, we denote the set of all short sequences as $\mathcal{D}_{ssl}$. To verify the accuracy of sequence purposes generated by the LLM, we manually label a portion ($20\%$) of short sequences as the ground truths. The LLM achieves the desirable accuracy of 92%.

### v2

# Harnessing Large Language Models for Group POI Recommendations

@@ Abstract.
The rapid proliferation of Location-Based Social Networks (LBSNs) has underscored the importance of Point-of-Interest (POI) recommendation systems in enhancing user experiences. While individual POI recommendation methods leverage users’ check-in histories to provide personalized suggestions, they struggle to address scenarios requiring group decision-making. Group POI recommendation systems aim to satisfy the collective preferences of multiple users, but existing approaches face two major challenges: diverse group preferences and extreme data sparsity in group check-in data. To overcome these challenges, we propose LLMGPR, a novel framework that leverages large language models (LLMs) for group POI recommendations. LLMGPR introduces semantic-enhanced POI tokens and incorporates rich contextual information to model the diverse and complex dynamics of group decision-making. To further enhance its capabilities, we developed a sequencing adapter using Quantized Low-Rank Adaptation (QLoRA), which aligns LLMs with group POI recommendation tasks. To address the issue of sparse group check-in data, LLMGPR employs an aggregation adapter that integrates individual representations into meaningful group representations. Additionally, a self-supervised learning (SSL) task is designed to predict the purposes of check-in sequences (e.g., business trips and family vacations), thereby enriching group representations with deeper semantic insights. Extensive experiments demonstrate the effectiveness of LLMGPR, showcasing its ability to significantly enhance the accuracy and robustness of group POI recommendations.
@@
copyright: nonejournalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Korea.booktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreaisbn: 979-8-4007-2040-6/2025/11doi: 10.1145/XXXXXX.XXXXXX
@@ 1. Introduction
With the rapid expansion of Location-Based Social Networks (LBSNs) such as Foursquare and Weeplace, Point-of-Interest (POI) recommendations have become essential for helping individuals discover places that align with their interests (Zhang et al., 2025; Yin et al., 2015; Huang et al., 2024a). These recommendations are crucial for applications like mobility prediction, route planning, and location-based advertising (Li et al., 2018; Yin and Cui, 2016; Yin et al., 2016). While individual POI recommendation systems effectively cater to personal preferences by utilizing users’ check-in histories, they fall short in scenarios requiring group decision-making. As location-based social activities become more prevalent, the demand for systems that address the preferences of multiple users simultaneously has transformed into a pressing need. This has given rise to the concept of Group POI Recommendation (Li et al., 2020; Liu et al., 2024), which is increasingly essential in scenarios such as friends planning a weekend outing, families on vacation, or colleagues organizing a business trip, where satisfying the collective preferences of all group members is crucial.
@@
Group POI recommendation systems typically involve a two-phase pipeline. In the first phase, the group is treated as a virtual user, allowing traditional individual POI recommendation methods, such as matrix factorization (Lian et al., 2014), transformer-based approaches (Luo et al., 2021), graph-based approaches (Wang et al., 2022), and diffusion-based approaches (Qin et al., 2023), to be directly applied. These methods learn group representations based on group-level check-in data and individual representations based on individual-level check-in data, respectively. In the second phase, these group representations are further enhanced by aggregating the individuals’ representations within the group. Aggregation approaches include the simple average (Vinh Tran et al., 2019), weighted average (Sojahrood and Taleai, 2021), attention mechanism (Sankar et al., 2020), and graph convolution network (Li and Zhu, 2024). However, despite their effectiveness, existing group POI recommendation systems predominantly rely on ID-based data, utilizing only sparse individual and group check-in data to capture statistical patterns, leaving the semantic richness inherent in POI text data untapped.
@@
Benefiting from the recent breakthroughs in pre-trained large language models (LLMs), which have significantly advanced the semantic understanding of natural language, researchers have begun exploring the integration of LLMs into POI recommendation systems. This integration aims to better capture the rich semantic information embedded in POI data (e.g., POI categories), thereby alleviating data sparsity and cold-start issues. For example, LLMMove (Feng et al., 2024) and LLM-Mob (Wang et al., 2023) utilize a series of carefully designed prompts to input individual users’ check-in records into pre-trained LLMs, leveraging the commonsense knowledge embedded in the LLMs to better exploit the rich semantic information in check-in data. Additionally, LLM4POI (Li et al., 2024) further fine-tunes the pre-trained LLM using individual check-in data to enhance its performance.
Although existing LLM-based POI recommendation methods have shown effectiveness, they are primarily designed for individual POI recommendations and face significant challenges when applied to the group POI recommendation scenario. These challenges arise from two major issues:
  • Diverse Preferences and Complex Group Decision-making. Unlike individual POI recommendations, which aim to satisfy the preferences of a single user (as illustrated in Figure 1), group POI recommendations require reaching a consensus among group members on a shared POI to visit. This is particularly challenging because the preferences of individuals within a group are often diverse, and, in some cases, the group’s overall preference may differ from those of its individual members (Guo et al., 2021). Modeling this intricate decision-making process is a complex task.
  • Extreme Data Sparsity. Group check-in data is inherently much sparser than individual check-in data, as it involves multiple users visiting the same POI simultaneously (e.g., a group of friends dining together at a restaurant). For example, as discussed in Section 4.1, group check-in data in three public datasets are approximately 50–100 times sparser than individual check-in data. While existing LLM-based POI recommendation methods have demonstrated potential in mitigating data sparsity at the individual level by utilizing semantic information, the extreme sparsity of group check-in data presents an even greater challenge.
@@
To address the challenges of group POI recommendation, we propose LLMGPR, a novel framework that leverages pre-trained large language models (LLMs) to empower Group POI recommendation. To tackle the diverse group preferences challenge, our framework first introduces semantic-enhanced POI tokens, which are integrated alongside the original word tokens of the LLM. The embeddings for these POI tokens are initialized by extracting rich textual information (e.g., category and address) using the LLM. By combining semantic-enhanced POI tokens with contextual information such as positional encodings, spatio-temporal differences, and sequential patterns, our model captures the multifaceted factors influencing group decision-making. These include sequential effects, geographical influences, temporal patterns, and semantic relationships (Xie et al., 2016). Building on this, we propose a sequencing adapter based on Quantized Low-Rank Adaptation (QLoRA), which processes the representations of check-in activities to generate check-in sequence representations. These representations can then be flexibly applied as either group-level or individual-level representations, depending on the sequence type.
@@
To address the extreme data sparsity challenge in group check-in data, we introduce two key components. First, an aggregation adapter combines individual member representations derived from their respective check-in sequences into a unified group representation. Second, we design a self-supervised learning (SSL) task to further enrich the group representations. This task predicts the underlying purposes of check-in sequences (e.g., business trips and family vacations) by utilizing spatio-temporal and contextual cues within each sequence. Specifically, we employ a purpose-oriented prompt to guide the pre-trained LLM in categorizing sequences, which then serves as ground truth for self-supervised learning. The SSL signal enhances the model’s ability to extract semantic information from sparse group-level check-in data, refining group representations.
By jointly optimizing the SSL task and the adapters for the recommendation task, LLMGPR effectively leverages the rich semantic and contextual information in check-in data, achieving robust and accurate group POI recommendations.
The primary contributions of this work are summarized as follows:
@@
  • We propose LLMGPR, the first pre-trained LLM-based framework for group POI recommendation.
@@
  • We introduce semantic-enhanced POI tokens and incorporate rich contextual information to model group check-in activities. Additionally, we propose a sequencing adapter based on QLoRA to learn group representations from check-in sequences.
  • We address the challenge of sparse group-level check-in data by developing an aggregation adapter to combine individual representations into group representations and introducing a SSL task that predicts check-in sequence purposes, enabling the model to extract meaningful semantic insights.
@@ 2.1. Notations
Definition 2: POI Token. As the extensions of the original word tokens, POI tokens encapsulate unique IDs and encoded representations corresponding to specific POIs. To enrich these tokens with semantic information, their embeddings are initialized by analyzing related information with the large language model, and the details are shown in Section 3.1.1. All POI embeddings are stored in the POI embedding matrix $\textbf{E}_{poi}$.
@@ 3. Methodology
3.1. Group POI Recommendation with LLM
@@
3.1.1. Learning Embeddings of POI Tokens
Before establishing the sequence representation semantically, we first introduce POI tokens and initial corresponding embeddings alongside the original word tokens of the LLM. While semantic representations capture the nuanced meanings of POI names, categories and other information, relying solely on semantics may not fully leverage the unique identifiers of POIs. By incorporating POI embeddings through their IDs, we can ensure that each location is distinctly recognized and appropriately contextualized. This method allows us to directly reference specific POIs, enhancing the precision of recommendations and maintaining the integrity of location-specific data. To balance the depth of semantic understanding with the specificity of location identifiers, the embedding of each POI is initialized by the large language model with a carefully designed prompt which is shown as below:
Generate an embedding for the following Point of Interest (POI).
Name: ¡name¿
Category: ¡category¿
 Description: ¡description¿
 Reviews: ¡reviews¿
 Latitude: ¡latitude¿
Longitude: ¡longitude¿
Address: ¡address¿
where we directly use the average of the final-layer token embeddings as the initialized POI embedding without going through the linear and softmax layers. Applying this process to all POIs results in an embedding matrix $\textbf{E}_{poi}\in\mathbb{R}^{P\times d}$, where $P$ is the number of all POIs and $d$ is the embedding size which is the same as the size of the token embedding. In addition, these embeddings will be further fine-tuned during the following group POI recommendation task with specific objectives.
3.1.2. Generating Candidate Group POI
Given all POI embeddings, we design another prompt shown as follows to generate the representation of the check-in sequence.
Generate an embedding for the provided check-in sequence, which has multiple check-in activities, and the format of a check-in activity is:
(POIID, timestamp, temporal difference from previous check-in, spatial difference from previous check-in).
The check-in sequence is: ¡check-in sequence¿
When this prompt is processed by the word embedding layer, all ”POIIDs” of check-in activities will be regarded as special POI tokens instead of standard word tokens. Therefore, they are encoded by querying the POI embedding matrix $\textbf{E}_{poi}$ with their IDs. These POI embeddings, along with the embeddings of other standard word tokens, are then fed into the subsequent layers of the large language model (LLM) for further processing. Then, the average of the final-layer token embeddings will be used as the embedded representation of the check-in sequence, denoted as:
(4) $\textbf{e}_{g_{n}}=AVG(\Theta(\mathcal{X}_{g_{n}},\textbf{E}_{poi})),$
@@ 3.1.2. Generating Candidate Group POI
3.1.3. Aligning LLM with Group POI Recommendation
@@
3.2. Enhancing Group Representations with Individual Check-ins
@@ 3.2.1. Aggregating Individual Rereprentations with LLM
3.2.2. Fine-tuning with QLoRA for Representation Refinement
@@
3.3. Addressing Data Sparsity with Self-supervised Leaning
As mentioned above, the LLM will be fine-tuned for check-in sequence representation generation and individual preference aggregation. However, the sparsity of user-POI interactions has been a long-standing issue in POI recommendation, which deteriorates further in group settings. To this end, we introduce a self-supervised signal to enhance representations for check-in sequences with the corresponding purposes, such as business trips and tourism, as the travel purposes provide a strong signal to predict the next POI a group/user will visit. Unfortunately, check-in data typically lacks explicit labels indicating the purpose of each check-in sequence. To overcome this limitation, we leverage the powerful capabilities of the LLM to classify each check-in sequence. The LLM analyzes the spatio-temporal and contextual information within the sequence and assigns a category that reflects its underlying purpose. To achieve this, we design the corresponding prompt which is shown as follows:
Given a sequence of check-in activities, your task is to classify into one of the following categories:
{Business Trip, Work Commute, Tourism, Short Vacation, Cultural & Entertainment, Shopping & Dining, Social Events, Fitness & Wellness, Daily Commute, Medical & Health Visits, Education & Training}.
The format of a check-in activity is: (POIID, Name, Category, longitude, latitude, temporal difference from previous check-in, spatial difference from previous check-in)
The check-in sequence is: ¡check-in sequence¿
The format of your response: This sequence is classified as [Category].
As shown in the above prompt, we have $11$ purpose labels: {Business Trip, Work Commute, Tourism, Short Vacation, Cultural & Entertainment, Shopping & Dining, Social Events, Fitness & Wellness, Daily Commute, Medical & Health Visits, Education & Training}. These labels are designed to reflect the most common and distinct purposes for short-term trips or sequences of check-ins. Each category is tied to a specific motivation or context that drives user mobility, ranging from daily routines to more occasional events. Together, they comprehensively capture the various reasons behind short-duration check-ins, ensuring that each sequence can be classified meaningfully and relevantly for POI prediction. Beyond that, those purpose labels are only valid for short sequences, since a purposeful trip won’t last long. Therefore, we divide the whole sequence into subsequences if the time difference between any two consecutive check-in activities is larger than five days. Here, we denote the set of all short sequences as $\mathcal{D}_{ssl}$. To verify the accuracy of sequence purposes generated by the LLM, we manually label a portion ($20\%$) of short sequences as the ground truths. The LLM achieves the desirable accuracy of 92%.